OpenStudy (jocelm):

Factor 4x²+9 I forgot how to do this, want someone to show me.

TheSmartOne (thesmartone):

\(\Large a^2 + b^2 =(a - bi)(a+bi)\) where i is the imaginary root, i = sqrt(-1)

OpenStudy (jocelm):

Oh wow, okay and is that like a formula you just used that can be used for other equations?

TheSmartOne (thesmartone):

There are a couple other formulas, mhmm \(\Large a^2 - b^2 = (a+b)(a-b)\) \(\Large a^2 + 2ab + b^2 = (a+b)^2\) \(\Large a^2 - 2ab + b^2 - (a-b)^2\) \(\Large a^3 - b^3 = (a - b)(a^2+ab+b^2)\) \(\Large a^3 + b^3 = (a+b)(a^2-ab+b^2)\)

OpenStudy (jocelm):

Thanks so much I couldn't remember!

TheSmartOne (thesmartone):

Eventually with practice, you should be able to memorize these formulas. They'll help you solve questions much faster (: